### v1

# RAG-ConfusionQA: A Benchmark for Evaluating LLMs on Confusing Questions

Abstract
Conversational AI agents use Retrieval Augmented Generation (RAG) to provide verifiable document-grounded responses to user inquiries. However, many natural questions do not have good answers: about 25% contain false assumptions Yu et al. (2023), and over 50% are ambiguous Min et al. (2020). RAG agents need high-quality data to improve their responses to confusing questions. This paper presents a novel synthetic data generation method to efficiently create a diverse set of context-grounded confusing questions from a given document corpus. We conduct an empirical comparative evaluation of several large language models as RAG agents to measure the accuracy of confusion detection and appropriate response generation. We contribute a benchmark dataset to the public domain.
RAG-ConfusionQA: A Benchmark for Evaluating LLMs on Confusing Questions
Zhiyuan Peng Santa Clara University Santa Clara, CA, USA zpeng@scu.edu                        Jinming Nian Santa Clara University Santa Clara, CA, USA jnian@scu.edu                        Alexandre Evfimievski aevfim@outlook.com                        Yi Fang Santa Clara University Santa Clara, CA, USA yfang@scu.edu
1 Introduction
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents Lewis et al. (2020); Guu et al. (2020); Gao et al. (2024). The RAG technology enables a large language model (LLM) to generate responses to users’ inquiries with information from a curated knowledge base stored as a collection of text documents. For each inquiry, a typical RAG pipeline retrieves a relevant document and uses it to generate an LLM response. Since an LLM is not a reliable source of knowledge due to its propensity to hallucinate, in business critical settings its role must be limited to language manipulation and common-sense reasoning. Every domain-specific claim in the response must be verifiable and traceable to its source document.
The performance of a RAG system is determined by evaluating its responses along a number of dimensions, including accuracy, faithfulness, completeness, answer relevance, and context relevance, as well as style metrics Es et al. (2024); Saad-Falcon et al. (2024); Liu et al. (2024); Zheng et al. (2023). Besides, the assumption that a user’s question is answerable may not hold. A well-designed RAG system should implement an exception-handling pathway that detects confusing or unanswerable questions and responds by pointing out the confusion or seeking help. Confusing questions in RAG are the subject of this work.
There is substantial prior work on using LLM to identify and respond to questions with a false premise, as well as using such questions as benchmarks for LLM evaluation Zhu et al. (2024); Yuan et al. (2024); Hu et al. (2023). Most of that work focuses on “common-sense” false premises that contradict basic knowledge learned by LLMs at pretraining. They do not consider premises that contradict a retrieved document, or appear to do so due to the lack of clarifying context such as timing. Our goal is to help LLMs respond to a variety of types of confusing questions, especially those confusing in the context of the document.
Building a “deconfusion” system involves addressing 3 challenges: creating a labeled dataset with many diverse examples of confusing questions; building a classifier to detect confusion; and, developing a proper response generator for each supported confusion type. As human content creation is expensive, we focus on LLM-based synthetic generation of confusing questions for a given set of documents. Our contributions are as follows:
  • We design, implement, and evaluate a novel synthetic data generation method to efficiently generate diverse high-quality confusing questions for a set of provided documents.
  • We discuss the types of confusing questions in RAG setting and the common situations where confusion arises.
  • We perform an empirical comparative evaluation of several LLMs as RAG agents to measure the accuracy of confusion detection and appropriate response generation.
  • We will release our code and the RAG-ConfusionQA dataset upon acceptance of the paper [^1] https://github.com/zhiyuanpeng/RAG-ConfusionQA.git.
2 Related Work
Benchmark Datasets
Most of the recent work on LLM response to confusing questions focuses on stand-alone questions to be answered based on the LLM’s pretraining knowledge, without a context document. The questions may be naturally collected from online sources Li et al. (2024); Yu et al. (2023), written by human annotators Hu et al. (2023), or synthetically generated Zhu et al. (2024); Yuan et al. (2024). Released benchmark datasets with confusing questions include KG-FPQ Zhu et al. (2024), FalseQA Hu et al. (2023), FLUB Li et al. (2024), and CREPE Yu et al. (2023). Ambiguous open-domain questions were studied in an earlier work Min et al. (2020) that released AmbigNQ, a dataset of natural ambiguous questions.
Question Generation
Zhu et al. (2024); Yuan et al. (2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of subject + relation + object using Wikipedia data, replacing the object with a similar, but wrong, object, then filling in a template to generate a false premise question. Content-grounded (non-confusing) synthetic data generation is studied in Yehudai et al. (2024); Wang et al. (2023b); Lee et al. (2023). Given the grounding passages, they generate data entries by few-shot prompting with task-specific examples, then filter to ensure data quality, faithfulness, and diversity.
False Premise Detection
by LLMs is commonly achieved with fine-tuning, chain-of-thought (CoT), and in-context learning Zhu et al. (2024); Li et al. (2024); Hu et al. (2023). Experiments show that fine-tuning outperforms the other two methods Hu et al. (2023). CoT does not bring stable improvements, while in-context learning can bring performance improvement if more examples are shown to the LLM Li et al. (2024). In Yu et al. (2023) a retrieval model finds relevant Wikipedia passages to send to a RoBERTa-based multi-passage classifier. Yuan et al. (2024) deep-dives into the LLM and finds $\sim$1of attention heads primarily responsible for getting confused by false premise questions. Constraining these attention heads during inference yields a performance increase of nearly 20%. They also observed a strong correlation between hallucination and response uncertainty metrics.
Context Based Verification
Natural questions unanswerable due to unverifiable premises were studied in Kim et al. (2021). This work is closer to RAG because it associates a document with each question: a Wikipedia page retrieved by the Google search engine. A premise is unverifiable (rather than false) if it not entailed by the document. The proposed system extracts all premises from a question, verifies them by checking entailment from each sentence of the document, then generates a response stating the unverifiable premise if found.
3 Methods
We built benchmark RAG-ConfusionQA to evaluate LLMs on the context-grounded response or answer generation step of the RAG pipeline when the questions are confusing (Section 3.1). In this step, given a standard prompt $p$ (Appendix F.6), a confusing question $q$, and a retrieved document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question, a process we refer to as defusion (or de-confusion), rather than attempting to answer the question directly. Document retrieval and ranking are out of scope for this work.
3.1 Confusing Questions
We call question $q$ confusing given document $d$ if $q$ cannot be answered faithfully and completely based on $d$ alone, without requesting — or providing — some clarification. Common confusion types in RAG-ConfusionQA (Appendix D) include:
False premise:
$q$ makes a wrong or a highly biased assumption that must be clarified. The contradiction may be with $d$ or with basic knowledge.
Not Mention:
Some in-domain information required to answer $q$ is not provided in $d$.
Ambiguous:
$q$ is missing some context or contains generic words with multiple interpretations, allowing different answers in the context of $d$.
Note that $q$ is interpreted within the context of $d$. If the name “John Smith” appears in both, it is assumed to refer to the same person unless stated otherwise. Similarly, $q$ and $d$ are assumed to refer to the same time, place, or use-case unless clarified, otherwise leading to confusion.
Algorithm 1 The guided hallucination method
1:Call LLMq to extract from $d$ a list of $n$ claims $c_{1},\ldots,c_{n}$; batching claims reduces LLMq-calls (Appendix F.1)
2:Partition $\{1,\ldots,n\}$ into disjoint subsets $S_{1},\ldots,S_{k}$, e.g. $S_{j}=\{i:i\mathop{\textrm{mod}}3=j-1\}$ for $j=1,2,3$
3:for several rounds (e.g. 3), and in each round, for all $S_{j}$ in partition $(S_{1},\ldots,S_{k})$ do
4:     In the claims list $c_{1},\ldots,c_{n}$ replace all claims $c_{i}$ where $i\in S_{j}$ by text “(missing)
5:     Provide LLMq with this modified list (but not $d$) and prompt it to recover the missing claims (Appendix F.2)
6:     In the claims list $c_{1},\ldots,c_{n}$ replace the missing claims with the imputed claims
7:end for
8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only novel claims (Appendix F.9)
9:Call LLMq to generate one short question per each novel claim, focusing on one key element of it (Appendix F.8)
10:Call LLMq to filter out any questions answerable in the context of $d$ (Appendix F.3)
3.2 Data Generation
Here we describe our method of generating confusing questions for a given set of documents. We require each document to be short enough to fit in the LLM prompt along with instructions and examples, but long enough to make at least 5 to 10 separate claims (i.e. a few paragraphs long). We prefer documents that are novel to the LLMs under evaluation, e.g. news articles published later than all pretraining cutoff dates, so that the LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Thus, we collected 200 news articles published in 2024, each containing more than 150 words for each topic (Appendix E). Sentences were selected sequentially from the top until the total word count exceeded 300, using the NewscatcherBugara et al. (2020). Given a set of such documents, we generate multiple confusing questions about each document, along with some non-confusing ones for comparison.
We start by picking LLMq (GPT-4o-mini by default in this paper), the “teacher” LLM that will generate the confusing questions. We prefer, but do not require, LLMq to be larger and stronger than the LLMr under evaluation. For each document $d$, we call LLMq to generate a list of factual claims, then iteratively replace subsets of claims with hallucinated claims while maintaining semantic similarity to $d$ using guided hallucination, see Algorithm 1. We also generate some non-confusing questions for each document $d$ by calling LLMq either directly.
3.3 Evaluation and Metrics
We first assessed the quality of RAG-ConfusionQA by sampling a subset of questions and manually verifying whether they were indeed confusing or not (Section 3.4). Next, we evaluated how well various LLMr models could defuse the confusing questions in RAG-ConfusionQA. Given the large number of confusing questions, manual verification of LLMr responses is not feasible. Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question. To validate the effectiveness of this defusion detection method, we compared its performance with human annotations on a sampled dataset (Section 3.4). Formally, we define our tasks as follows:
Question generation:
What is the quality of confusing and non-confusing questions generated by Algorithm 1? (Section 3.4)
Defusion detection:
How accurately can AutoDefuseEval detect if an LLMr response to a confusing question defuses the confusion? (Section 3.4)
Confusion response:
How frequently does an LLMr, when responding to a confusing question using the standard prompt, generate a response that successfully defuses the confusion? (Section 4.1)
Confusion detection:
How accurately can LLMs identify which context-grounded questions are confusing and require special handling? (Section 4.2)
We evaluate “question generation” and “defuse detection” through accuracy and confusion matrix. We only apply accuracy for “confusion response” and “confusion detection”. Algorithm A has the steps we run for each LLMr and for each $(q,d)$ pair. We run steps 1, 2, and 3 for “confusion detection” and 1, 4, 5, and 6 for “confusion response”.
Following the self-consistency method Wang et al. (2023a), we perform multiple LLM calls in steps 2 and 5 of Algorithm A, taking the majority vote to determine the final label. We also present the LLMq with few-shot exemplars of questions, responses, and explanations (Section 4.3).
Model business entm food music news politics science sport tech travel Avg Std Dev
GPT-3.5 18.90 12.39 13.55 15.48 15.15 17.13 10.81 $20.24^{\star}$ 17.12 10.01 15.08 3.21
Llama 3.2 3B 68.34 66.11 64.89 67.90 67.50 66.94 57.98 $76.37^{\star}$ 67.07 60.06 66.31 4.70
Mistral 7B v0.3 71.27 70.37 65.81 71.60 71.62 69.26 60.35 $79.00^{\star}$ 69.97 64.29 69.35 4.77
Llama 3.1 70B 74.48 72.55 70.33 75.31 71.33 71.39 65.60 $\textbf{83.33}^{\star}$ 73.17 70.38 72.79 4.34
Llama 3.1 8B 76.18 74.03 70.23 76.07 73.15 72.87 64.37 $81.45^{\star}$ 73.47 68.11 72.99 4.45
Table 1: Evaluation of LLMs’ accuracy in defusing confusing questions from the RAG-ConfuseQA-Silver dataset across diverse news topics. For each LLM, the underscored value and stared value are the minimum and maximum values, respectively, across all the topics. Bold values are the maximum values across all the LLMs.
3.4 Human Annotation
To evaluate “question generation” and “defuse detection” (Section 3.3), we recruited six computer science graduate students for manual labeling (Appendix B). As reflected in “Ground Truth Acc,” of Appendix B, Table  3, our method achieved 94.91% accuracy in generating confusing and non-confusing questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion. As shown in Appendix B, Figure 2, the confusion matrix on the left strongly agrees with the ground truth, with minimal errors (5 false negatives and 6 false positives). The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
4 Experimental Results
4.1 Confusion Response
We selected several open-source LLMs (Appendix 4) to evaluate their ability to generate responses that effectively defuse confusing questions with a standard prompt (Appendix F.6). As shown in Table 1, a) GPT-3.5 performs the worst among all six LLMs but with the smallest Std Dev across all topics, indicating consistent performance. On the other hand, Llama 3.1 70B and Llama 3.1 8B achieve comparable top performances with similar Std Dev, despite their differences in model size; b) most LLMs perform the worst in the “science” topic, except for GPT-3.5, for which “science” is the second worst. This is expected, as science topics often require more specialized domain knowledge to defuse confusion effectively; c) all LLMs achieve their best performance on the “sport” topic, demonstrating a clear trend. The strong performance across LLMs on the “sport” topic might be attributed to the relative simplicity and unambiguous nature of sports-related content.
4.2 Confusion Detection
By default, we use GPT-4o-mini as our LLMq for the task of “confusion detection,” which identifies whether the generated questions are confusing. In this section, we compare GPT-4o-mini with other LLMs. As illustrated in Figure 1, GPT-4o-mini achieves the best performance, while the smallest model, Llama 3.2 3B, performs the worst.
Llama 3.2 3BMistral 7BLlama 3.1 8Bgpt-3.5Llama 3.1 70Bgpt-4o-mini606060608080808010010010010079.6379.6379.6379.6385.1985.1985.1985.1986.5786.5786.5786.5787.9687.9687.9687.9693.5293.5293.5293.5294.9194.9194.9194.91Different LLMsAccuracy (%)
Figure 1: Evaluation of instruction-finetuned LLMs’ accuracy in classifying questions as confusing or not on the RAG-ConfusionQA-Gold dataset
Module Accuracy
Base (gpt-4o-mini) 87.61
+ examples 97.35
+ self-consistency (n=3) 98.23
+ self-consistency (n=9) 98.23
Table 2: Evaluation of different prompts in the task “defusion detection” on RAG-ConfusionQA-Golden.
4.3 Ablation Study
We examined the impact of various prompts on the task of “defusion detection.” As seen in Table 2, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $n=3$ further boosted performance to 98.23%. However, increasing the number of generated sequences beyond three did not lead to additional improvements. Since RAG-ConfusionQA-Golden is a smaller dataset, using $n=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger RAG-ConfusionQA-Silver dataset. Therefore, we adopted $n=9$ in our experiments to balance accuracy and expenses (Appendix F.10).
5 Conclusion
We introduced RAG-ConfusionQA, a benchmark designed to assess how effectively an LLM defuses confusion when responding to confusing questions using a standard prompt. Through human annotation, we demonstrated that RAG-ConfusionQA achieves a 94.91% accuracy in generating confusing and non-confusing questions, while our proposed AutoDefusionEval achieves a 98.23% accuracy in detecting whether an LLM’s response defuses confusion. Additionally, our evaluation highlights that current LLMs often attempt to answer confusing questions to varying degrees. For future work, we aim to develop methods to better detect confusing questions and address the tendency of LLMs to respond to them directly instead of defusing the confusion.
6 Acknowledgement
We would like to express our sincere gratitude to Xiaoxiao Shang, Ethan Lin, Wei Mo, Zhan Shi, Xinze Ye, and Haisong Wang for their invaluable contributions to data labeling.
7 Limitations
In this paper, we only evaluated LLMs’ abilities to generate responses that effectively defuse confusing questions when prompted with a standard prompt (Appendix F.6). We chose this prompt because, ideally, LLMs should be able to address confusion with a basic, straightforward prompt. However, we have not yet assessed the performance of LLMs across different prompt variations.
Appendix A Algorithm
Algorithm 2 Evaluation steps
1:Call LLMr on $(d,q)$ with the standard RAG prompt and get its response $r$ (Appendix F.6)
2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is confusing given $d$ (Appendix F.3)
3:Aggregate $m$ predictions by majority vote
4:if $q$ is generated as confusing then
5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix F.4)
6:     Aggregate $m$ predictions by majority vote
7:end if
Appendix B Human Annotation
Figure 2: Confusion matrix of confusion and defusion on RAG-ConfusionQA-Golden.
Group Name Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
confusion defusion confusion defusion confusion defusion confusion defusion
1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
  — 87.50 94.59
2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
  — 91.67 90.00
3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
  — 98.61 100.00
Table 3: Agreement scores and the performance of our method on the RAG-ConfusionQA-Golden.
Following the methodology from Cochran (1977), representing approximately 20,000 synthetic questions in RAG-ConfuseQA at a 95% confidence level, with a 4% margin of error, and an anticipated accuracy of $90\pm 5\%$, required labeling at least 202 questions. We sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators. As shown in Table 3, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is confusing and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency. Two more annotators manually resolved all disagreements to establish ground truth labels. Six annotators are computer science graduate students recruited via email and incentivized with complimentary meals. Another two annotators are the authors of this paper, who are also majoring in computer science.
We used two metrics to assess our method: “Annotator Acc” and “Group’s Agree Acc.” “Annotator Acc” treats each annotator’s labels as ground truth, computing the accuracy of our method’s generated labels. “Group’s Agree Acc” uses the instances where both annotators within the same group agreed, treating this consensus as the ground truth. This method demonstrated better accuracies than “Annotator Acc,” leading us to resolve disagreements with two additional annotators, ultimately establishing the final labeled data as ground truth.
Appendix C Different LLMr
Model Name Quantization Knowledge Cut-off Date
Llama 3.2 3B Instruct Turbo FP8 2023/12
Llama 3.1 8B Instruct Turbo FP8 2023/12
Llama 3.1 70B Instruct Turbo FP8 2023/12
Mistral (7B) Instruct v0.3 FP16 2024/5
gpt-3.5-turbo   — 2021/9
Table 4: Different LLMr to be evaluated. In RAG-ConfuseQA-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cut-off date 5/22/2024.
Appendix D Confusing Types
Document:
Justin Jefferson, CeeDee Lamb, NFL Injury Statuses and Fantasy Impact for Week 3 Stephen Maturen/Getty Images
Two of the best wide receivers in the NFL are expected to be on the field for their respective Week 3 games. Justin Jefferson and CeeDee Lamb each received positive prognosis about their injury issues, which is something that can’t be said for the rest of the stars across the NFL. Christian McCaffrey is already on injured reserve, Deebo Samuel is out for a weeks and a slew of other running backs and wide receivers are dealing with ailments that could keep them out of Week 3. Below is a look at all of the significant injuries that could affect fantasy football matchups across Week 3. Justin Jefferson Off Injury Report
Justin Jefferson was taken off the Minnesota Vikings injury report on Friday. Jefferson’s status was up in the air because of a quad injury, but he practiced well enough this week that the injury is not a concern. The superstar wide out will be needed for Minnesota’s home clash with the Houston Texans, which has the potential to be a high-scoring affair. Jefferson is always the primary target in Minnesota when healthy, but he should have more targets in Week 3 because Jordan Addison and T.J. Hockenson are still out. Jefferson earned seven targets from Sam Darnold in Week 2. Only running back Aaron Jones had more than four targets against the New York Giants. Houston’s defense may give Jefferson some fits, led by cornerback Derek Stingley Jr. but it has allowed 330 receiving yards to opposing wide outs on just 20 catches through two weeks. Jefferson is an automatic start whenever he’s healthy, and his star power may be needed more on certain fantasy rosters in Week 3 depending on how many injuries affect a single roster.
Question 1: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
This question is confusing as it asks about "Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft" which was never discussed in the document. So this belongs to Not Mention
Question 2: How will Justin Jefferson’s struggles with hamstring issues in Week 3 affect the Vikings’ chances of winning against the Houston Texans?
This question assumes that Justin Jefferson is dealing with hamstring issues, which is not mentioned in the document. The document only refers to a quad injury that Jefferson had, which is no longer a concern as of Week 3. The document makes it clear that Jefferson is healthy and will play without any restrictions. Therefore, the question is based on a False Premise
Question 3: How will his performance impact the team’s chances this season, considering they are missing key players?
This question uses the Ambiguous pronoun “his” without specifying who it refers to, making it unclear whether the question is asking about Justin Jefferson, CeeDee Lamb, or another player mentioned in the document. Additionally, “the team” is not defined explicitly, as both the Minnesota Vikings and other teams are referenced in the document. Without clarifying which player or team the question is referring to, it is impossible to provide an accurate answer based on the document.
Appendix E Statistics of RAG-ConfusionQA
Attr Golden Silver
# of documents 45 2,000
# of non-confusing questions 103 9,052
# of confusing questions 113 10,218
average words per document 280 276
average words per non-confusing question 16 16
average words per confusing question 16 16
Table 5: Statistics of RAG-ConfusionQA. Golden is sampled from Silver and annotated by annotators (Appendix B).
Appendix F Prompt
F.1 Extract Claims Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and list {num_fact} most important facts it contains. Each fact should be
stated in a clear, standalone sentence with sufficient context to be understood
independently, avoiding undefined pronouns. Ensure that each fact is directly derived from
the document and does not include any information not mentioned within it.
Document:
    """{document}"""
{num_fact} most important facts:"
F.2 Recover Missing Claims Prompt
{mdframed}
Read the document below with a list of {num_fact} facts it contains. Note that some of the
facts are missing represented by (missing). Your task is to guess the missing facts could
have said and complete the missing facts. Each fact should be stated in a clear, standalone
sentence with sufficient context to be understood independently, avoiding undefined pronouns.
Please complete the missing facts and return all the {num_fact} facts in original order. You
must return all the {num_fact} facts.
Document:
    """{document}"""
Completed list of facts:"
F.3 Confusion Judgement Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and the question. Check if the question contains any confusing
parts, false assumptions, incorrect information, or mentions of people, locations,
organizations, or other entities that are not present in the document.
Document:
    """{document}"""
Question:
    {question}
Does the question contain any confusing parts, false assumptions, incorrect
information, or mentions of people, locations, organizations, or other entities
that are not present in the document? Please start by explaining your reasoning.
End your response with ’The answer is: Yes or No.’:
F.4 Defusion Judgement Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
Document:
    """{example_document}"""
Question:
    {example_question}
{defuse_answer_examples}
{Defusion Instruction}
{defuse_reasoning_examples}
Read the document and answer the question based on the document.
Document:
    """{document}"""
Question:
    {question}
{response_from_gpt3.5}
{Defusion Instruction}
F.5 Defusion Instruction
{mdframed}
The question presented is confusing and unanswerable based on the document provided.
Read the response carefully. If the response attempts to answer the question,
like answering ’Yes’ to agree the question or ’No’ to disagree the question,
you should immediately conclude with The answer is: No. even if it later
acknowledges the questions confusion. Otherwise, determine if the response
clarifies that the question cannot be answered or confusing or matches any of the
issues listed:
    False Premise: The response points out that the question is based
    on an incorrect assumption.
    Not Mentioned: The response indicates that the information needed to answer
    the question is not available or not mentioned in the document.
    Ambiguous: The response points out unclear pronouns or vague terms in the
    question that prevent a clear answer.
Begin with your reasoning and finally conclude with The answer is: Yes. or
The answer is: No.. Note that if the response attempts to answer the question
at any point, like answering ’Yes’ to agree the question or ’No’ to disagree the
question, you should immediately conclude with The answer is: No. even if it
later acknowledges the questions confusion:
F.6 RAG Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
Document:
    """{document}"""
Question:
    {question}
F.7 Non-confuing Question Generation Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document attentively and compile a numbered list of the top {num_q} questions
that the document directly answers. Ensure each question is clear, accurate, and devoid
of confusion, false assumptions, undefined pronouns, or misinformation. Avoid referencing
people, locations, organizations, or other entities not explicitly mentioned in the
document. Construct each question to be thought-provoking, containing between 13 to 18
words, and sufficiently detailed to avoid being overly straightforward.
Document:
    """{document}"""
Questions:
F.8 Confusing Question Generatioin Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and review the list of hallucinated facts. For each hallucinated fact,
craft a single, specific and concise question containing 13 to 18 words that incorporates
the key element of the fact, ensuring the question is intentionally confusing.
The question should not be answerable using any information present in the document.
The question should not combine multiple queries and each question should address only
one specific aspect. If a question cannot be formulated for a particular hallucinated
fact, you may omit it.
Document:
    """{document}"""
hallucinated facts:
    {hallucinated_facts}
Questions:
F.9 Remove Claims Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document below with a list of {num_true_fact} ground-truth facts it contains
and a list of {num_false_fact} hallucinated facts that are not supported by the document.
Your task is to remove any hallucinated facts that can be supported by either the
document or the {num_true_fact} ground-truth facts.
Please only return the remaining hallucinated facts, along with their original order numbers.
Document:
    """{example_document}"""
4 ground-truth facts:
    {true_facts}
{num_false_fact} hallucinated facts:
    {false_facts}
Remaining hallucinated facts:
    {remained_facts}
Read the document below with a list of {num_true_fact} ground-truth facts it contains
and a list of {num_false_fact} hallucinated facts that are not supported by the document.
Your task is to remove any hallucinated facts that can be supported by either the
document or the {num_true_fact} ground-truth facts.
Please only return the remaining hallucinated facts, along with their original order numbers.
Document:
    """{document}"""
{num_true_fact} ground-truth facts:
    {true_facts}
{num_false_fact} hallucinated facts:
    {hallucinated_facts}
Remaining hallucinated facts:
F.10 Expenses
Generating the RAG-ConfusionQA dataset using GPT-4o-mini costs approximately \$50. For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/. Excluding the dataset generation, all other experiments cost around \$60.