### v2

# ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG

Abstract
Conversational AI agents use Retrieval Augmented Generation (RAG) to provide verifiable document-grounded responses to user inquiries. However, many natural questions do not have good answers: about 25% contain false assumptions Yu et al. (2023), and over 50% are ambiguous Min et al. (2020). RAG agents need high-quality data to improve their responses to confusing questions. This paper presents a novel guided hallucination-based method to efficiently generate a diverse set of borderline out-of-scope confusing questions for a given document corpus. We conduct an empirical comparative evaluation of several large language models as RAG agents to measure the accuracy of confusion detection and appropriate response generation. We contribute a benchmark dataset to the public domain.
ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG
Zhiyuan Pengthanks: Equal contribution. Santa Clara University Santa Clara, CA, USA zpeng@scu.edu                        Jinming Nian11footnotemark: 1 Santa Clara University Santa Clara, CA, USA jnian@scu.edu                        Alexandre Evfimievski aevfim@outlook.com                        Yi Fang Santa Clara University Santa Clara, CA, USA yfang@scu.edu
1 Introduction
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents Lewis et al. (2020); Guu et al. (2020); Gao et al. (2024). Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and provides it to a large language model (LLM) that generates a response. RAG performance is evaluated on dimensions like the response accuracy, faithfulness, completeness, answer and context relevance, and style Es et al. (2024); Saad-Falcon et al. (2024); Liu et al. (2024); Zheng et al. (2023). Errors may stem from a document-to-question mismatch, hallucinating parts of the response, wrong understanding of the document or the question.
Many user questions have no direct answer: some may rely on a false premise, some may be ambiguous in what they ask, and some are out of scope for the knowledge base. A well-designed RAG system should detect such confusing questions and respond appropriately, either by clarifying the confusion or seeking help.
Dataset RAG Setting Source Category Cut-off-knowledge Annotation Method
AMBIGQA Min et al. (2020) Wiki Multi-answers Human
CREPE Yu et al. (2023) Reddit False Premise Human
FalseQA Hu et al. (2023)  False Premise Human
KG-FPQ Zhu et al. (2024) Wiki False Premise LLM
FAITH Yuan et al. (2024) Wiki False Premise LLM
FLUB Li et al. (2024a) Baidu Tieba Cunning Texts Human
ScopeQA News Out-of-scope Human+LLM
Table 1: ScopeQA vs Existing Datasets
Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type. As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge but neglects out-of-scope queries, which lack answers in retrieved documents. Borderline out-of-scope questions can confuse LLMs into making up a hallucinated answer. This type of confusion is the focus of this paper.
While queries in MS MARCO Nguyen et al. (2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cut-off. ScopeQA fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to minimize human annotation effort. Our contributions are as follows:
  • We propose and implement a method to generate borderline out-of-scope questions for given documents efficiently.
  • We validate our synthetic data by training out-of-scope question detectors.
  • We evaluate LLMs as RAG agents for detecting confusion and generating responses.
  • Our code and dataset are released at https://github.com/zhiyuanpeng/ScopeQA.git.
2 Related Work
Benchmark Datasets
Most of the recent work on LLM responses to confusing questions focuses on stand-alone questions answered using LLM’s parametric knowledge without a context document. The questions may be naturally collected from online sources Li et al. (2024b); Yu et al. (2023), written by human annotators Hu et al. (2023), or synthetically generated Zhu et al. (2024); Yuan et al. (2024). A comparison of these released benchmark datasets is shown in Table 1.
Question Generation
Zhu et al. (2024); Yuan et al. (2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of subject + relation + object from Wikipedia, replacing the object with a similar but incorrect one, and filling in a template. Content-grounded (non-confusing) synthetic data generation is studied in Yehudai et al. (2024); Wang et al. (2023b); Lee et al. (2023). Given the grounding passages, they generate data entries by few-shot prompting with task-specific examples, then filter to ensure data quality, faithfulness, and diversity.
Mitigation
is commonly achieved with in-context learning  Zhu et al. (2024); Li et al. (2024b), chain-of-thought (CoT)  Wei et al. (2022) and fine-tuning. Experiments show that fine-tuning outperforms the other two methods Hu et al. (2023). CoT yields inconsistent improvements, while in-context learning improves performance with more examples Li et al. (2024b). Yuan et al. (2024) identifies  1% of attention heads responsible for confusion by false-premise questions, showing a 20% performance gain by constraining these heads.  Kim et al. (2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.
3 Methods
We developed the ScopeQA data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions. We assume that the retrieved document is the best available match for the question and focus on the response generation step of the RAG pipeline. In this step, given a standard prompt $p$, an out-of-scope question $q$, and a retrieved document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question, a process we refer to as defusion (or de-confusion), rather than attempting to answer the question directly and hallucinate.
Algorithm 1 The guided hallucination method
1:Call LLMq to extract from $d$ a list of $n$ claims $c_{1},\ldots,c_{n}$; batching claims reduces LLMq-calls (Appendix F.1)
2:Partition $\{c_{1},\ldots,c_{n}\}$ into disjoint subsets $S_{1},\ldots,S_{k}$, e.g. $S_{j}=\{c_{i}:i\mathop{\textrm{mod}}3=j-1\}$ for $j=1,2,3$
3:for several rounds (e.g. 3), and in each round, for all $S_{j}$ in partition $(S_{1},\ldots,S_{k})$ do
4:     In the claims list $c_{1},\ldots,c_{n}$ replace all claims $c_{i}$ where $i\in S_{j}$ by text “(missing)
5:     Provide LLMq with this modified list (but not $d$) and prompt it to recover the missing claims (Appendix F.2)
6:     In the claims list $c_{1},\ldots,c_{n}$, replace the missing claims with the claims recovered in above step.
7:end for
8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only out-of-scope claims (Appendix F.11)
9:Call LLMq to generate one short question per each novel claim, focusing on one key element of it (Appendix F.10)
10:Call LLMq to filter out any questions answerable in the context of $d$ (Appendix F.3)
3.1 Data Generation
Here we describe our method of generating out-of-scope questions for a given set of documents. We require each document to be short enough to fit in the LLM prompt along with instructions and examples but long enough to make at least 4 to 10 separate claims (i.e. a few paragraphs long). We prefer documents that are novel to the LLMs under evaluation, e.g. news articles published later than all pretraining cutoff dates, so that the LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Thus, by using the Newscatcher Bugara et al. (2020), we collected 200 news articles published in 2024, each containing more than 150 words, for each topic (Appendix E). Sentences were selected sequentially from the top until the word count exceeded 300. Given a set of such documents, we generate multiple out-of-scope questions about each document and some in-scope ones for comparison.
We start by picking LLMq (GPT-4o-mini by default in this paper), the “teacher” LLM that will generate the out-of-scope questions. We prefer but do not require LLMq to be larger and stronger than the LLMr under evaluation. For each document $d$, we call LLMq to generate a list of factual claims, then iteratively replace subsets of claims with hallucinated claims while maintaining semantic similarity to $d$ using guided hallucination, see Algorithm 1. We also generate in-scope questions for each document $d$ by calling LLMq directly.
3.2 Evaluation
We first assessed the quality of ScopeQA by sampling a subset of questions and manually verifying whether they were indeed out-of-scope or not (Section 3.3). Next, we evaluated how well various LLMr models could defuse the out-of-scope questions in ScopeQA. Given the large number of out-of-scope questions, manual verification of LLMr responses is not feasible. Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question. To validate the effectiveness of this defusion detection method, we compared its performance with human annotations on a sampled dataset (Section 3.3). Formally, we define our tasks as follows:
Question generation:
What is the quality of out-of-scope and in-scope questions generated by Algorithm 1? (Section 3.3)
Defusion detection:
How accurately can AutoDefuseEval detect if an LLMr response to a out-of-scope question defuses the confusion? (Section 3.3)
Out-of-scope response:
How frequently does an LLMr, when responding to an out-of-scope question using different prompting methods, generate a response that successfully defuses the confusion? (Section 4.1)
Out-of-scope detection:
How accurately can LLMs identify which context-grounded questions are out-of-scope and require special handling? (Section 4.2)
Appendix A has the steps we run for evaluating the above tasks. We run steps 1, 2, and 3 for “out-of-scope detection” and 1, 4, 5, 6, and 7 for “out-of-scope response”. Following the self-consistency method Wang et al. (2023a), we perform multiple LLM calls in steps  2 and  5 of Algorithm A, taking the majority vote to determine the final label.
Prompt Model business entm food music news politics science sport tech travel Avg Std Dev
Basic GPT-3.5 18.06 10.96 12.32 14.27 13.59 15.95 10.15 $18.37^{\star}$ 15.59 9.63 13.89 2.97
Llama 3.2 3B 63.59 61.55 61.09 62.07 63.49 64.01 53.35 $71.71^{\star}$ 63.69 56.51 62.11 4.60
Mistral 7B v0.3 68.40 67.63 60.88 67.74 67.96 66.86 56.07 $76.23^{\star}$ 66.87 61.06 65.97 5.20
Llama 3.1 8B 71.25 69.42 67.56 72.04 67.16 71.89 59.73 $77.90^{\star}$ 70.15 63.49 69.06 4.74
Llama 3.1 70B 71.54 70.22 67.56 71.07 67.66 70.75 61.51 $79.76^{\star}$ 70.67 68.04 69.88 4.33
Llama 3.3 70B 71.25 72.71 68.17 73.51 69.84 69.90 60.04 $\textbf{82.42}^{\star}$ 73.23 68.78 70.98 5.29
Two-shot GPT-3.5 65.65 65.24 57.60 65.88 62.20 66.48 52.41 $71.71^{\star}$ 59.90 56.83 62.39 5.43
Llama 3.2 3B 80.77 79.28 76.08 79.47 76.79 78.25 72.28 $84.09^{\star}$ 79.90 71.64 77.85 3.61
Mistral 7B v0.3 80.86 77.69 74.54 78.79 79.17 79.11 67.47 $86.84^{\star}$ 78.15 71.43 77.40 5.02
Llama 3.1 8B 83.32 80.98 80.18 84.16 78.77 82.53 72.70 $87.92^{\star}$ 82.77 78.41 81.17 3.87
Llama 3.1 70B 85.18 83.37 80.60 85.83 79.76 82.24 75.84 $\textbf{90.57}^{\star}$ 83.08 79.79 82.62 3.84
Llama 3.3 70B 85.48 81.27 80.39 84.26 79.27 82.34 75.84 $88.90^{\star}$ 83.49 78.62 81.99 3.57
Zero-shot-CoT GPT-3.5 63.00 61.95 55.95 64.03 62.40 62.58 53.03 $71.71^{\star}$ 59.69 56.30 61.07 4.96
Llama 3.2 3B 83.91 78.09 77.21 79.96 78.77 81.77 73.85 $85.46^{\star}$ 80.41 75.77 79.52 3.39
Mistral 7B v0.3 77.72 78.39 70.94 78.59 79.07 77.02 67.26 $84.77^{\star}$ 76.41 72.06 76.22 4.69
Llama 3.1 8B 79.39 78.09 74.85 78.40 78.67 79.20 70.29 $83.89^{\star}$ 77.64 74.50 77.49 3.43
Llama 3.1 70B 90.28 87.85 87.27 88.56 86.51 89.08 83.68 $91.55^{\star}$ 86.97 85.19 87.69 2.21
Llama 3.3 70B 93.52 91.24 90.55 92.57 90.77 92.12 89.12 $\textbf{95.09}^{\star}$ 93.85 90.90 91.97 1.71
Table 2: Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ScopeQA-Silver dataset across diverse news topics. For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics, within the same prompting method. Bold values are the maximum values for each topic across all the LLMs within the same prompting method.
3.3 Human Annotation
To evaluate “question generation” and “defusion detection” (Section 3.2), we recruited six computer science graduate students for manual labeling (Appendix B). As reflected in “Ground Truth Acc,” of Appendix B, Table  4, our method achieved 94.91% accuracy in generating out-of-scope and in-scope questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion. As shown in Appendix B, Figure 2, the confusion matrix on the left strongly agrees with the ground truth, with minimal errors (5 false negatives and 6 false positives). The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
4 Experimental Results
4.1 Out-of-Scope Response
We evaluated open-source LLMs (Appendix 5) on their ability to defuse out-of-scope questions with three prompting methods (Appendix F.6,  F.7,  F.8). As shown in Table 2, a) Two-shot prompt boosts all evaluated LLMs’ accuracy substantially, especially GPT-3.5, likely due to the examples clarifying how to respond to out-of-scope questions. In contrast, Zero-shot-CoT Kojima et al. (2022) merely invites the LLM to “reason step by step” beats Two-shot on three out of six LLMs. b) Larger models (70B) benefit more from Zero-shot-CoT than from Two-shot prompting, achieving accuracy gains of 5.07% to 9.98%, while smaller models (3B to 8B) experience only marginal changes, ranging from -3.68% to 1.67%. This difference likely stems from larger models’ ability to better utilize their advanced reasoning capabilities and extensive knowledge; c) Most LLMs perform worst on the “science” topic and best on “sport”. This is expected, as scientific content tends to contain more implicit premises assumed by the author rather than explicitly stated facts, thus requiring more domain knowledge. In contrast, sports content is simpler and unambiguous, and the facts in the document are often clearly stated.
4.2 Out-of-Scope Detection
By default, we use GPT-4o-mini as our LLMq for “out-of-scope detection” task. In this section, we study the impact of different LLMs and methods. As illustrated in Figure 1, “Direct Generate” prompts the LLM to do the detection (Appendix F.3); “Unused-token Classification” represents a classifier trained on ScopeQA-Silver using the embedding of an appended unused-token (Appendix H) as a feature. Finally, “SBERT NLI” represents a classifier trained on ScopeQA-Silver using embeddings from a sentence-transformer model (Appendix H.3). Among these methods, GPT-4o-mini performs best with the “Direct Generate” approach. Notably, ScopeQA proves its utility by enabling even smaller models, like the Llama 3.2 3B-based classifier, to achieve accuracy comparable to much larger models, such as Llama 3.1 70B and GPT-4o-mini. To minimize interference with the LLM’s generation process, we append an unused-token, leaving its potential impact for future investigation.
Figure 1: Evaluation of out-of-scope detection on ScopeQA-Gold.
Module Accuracy
Base (gpt-4o-mini) 87.61
+ examples 97.35
+ self-consistency (m=3) 98.23
+ self-consistency (m=9) 98.23
Table 3: Evaluation of different prompts for “defusion detection” on ScopeQA-Gold.
4.3 Ablation Study
We examined the impact of various prompts on “defusion detection” task. As seen in Table 3, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $m=3$ further boosted performance to 98.23%. However, further increasing $m$ did not lead to additional improvements. Since ScopeQA-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ScopeQA-Silver dataset. Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix G).
5 Conclusion
We introduced ScopeQA, a benchmark for evaluating LLMs’ ability to defuse out-of-scope questions in RAG. Human annotation validates that ScopeQA achieves 94.91% accuracy in distinguishing out-of-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses. We further demonstrated ScopeQA’s utility by training out-of-scope detectors. For future work, we plan to investigate the impact of appending an unused-token on the generation process. Additionally, we aim to explore instruction fine-tuning by training LLMs on ScopeQA and evaluating their performance. We also intend to apply for OpenAI funding to leverage more powerful models, such as GPT-4o, to enhance data quality.
6 Acknowledgement
We would like to express our sincere gratitude to Xiaoxiao Shang, Ethan Lin, Wei Mo, Zhan Shi, Xinze Ye, and Haisong Wang for their invaluable contributions to data labeling.
7 Limitations
In this paper, we evaluated LLMs’ abilities to defuse out-of-scope questions using three prompting methods: basic prompt, Two-shot prompt, and Zero-shot-CoT prompt. These methods were chosen because LLMs should ideally handle out-of-scope questions with straightforward prompts. However, we did not assess the performance of LLMs fine-tuned on ScopeQA-Silver due to computational resource constraints. Additionally, we empirically appended an unused-token to the end of the prompt for classification while minimizing its influence on LLMs’ generation processes. The effect of appending such tokens on generation remains unexplored and is left for future work. Finally, budget constraints prevented us from using more powerful models, such as GPT-4o or GPT-o1, which could theoretically produce higher-quality datasets. Future users can adopt our framework with more advanced models and alternative data sources to generate even higher-quality data tailored to their needs.
Appendix A Algorithm
Algorithm 2 Evaluation steps
1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix F.6,  F.7,  F.8)
2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix F.3)
3:Aggregate $m$ predictions by majority vote
4:if $q$ is generated as out-of-scope then
5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix F.4)
6:     Aggregate $m$ predictions by majority vote
7:end if
Appendix B Human Annotation
Figure 2: Confusion matrix of out-of-scope and defusion on ScopeQA-Gold.
Group Name Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
Out-of-Scope defusion Out-of-Scope Defusion Out-of-Scope Defusion Out-of-Scope Defusion
1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
  — 87.50 94.59
2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
  — 91.67 90.00
3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
  — 98.61 100.00
Table 4: Agreement scores and the performance of our method on the ScopeQA-Gold.
Following the methodology from Cochran (1977), we sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators. As shown in Table 4, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is out-of-scope and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency. Two more annotators manually resolved all disagreements to establish ground truth labels. All the annotators are computer science graduate students recruited via email and incentivized with complimentary meals. All annotators were trained by first reviewing the annotation guidelines and then labeling a small dataset previously labeled by the authors. The annotation guideline and annotation tool will be released along with the data and code upon acceptance.
We used two metrics to assess our method: “Annotator Acc” and “Group’s Agree Acc.” “Annotator Acc” treats each annotator’s labels as ground truth, computing the accuracy of our method’s generated labels. “Group’s Agree Acc” uses the instances where both annotators within the same group agreed, treating this consensus as the ground truth. This method demonstrated better accuracies than “Annotator Acc,” leading us to resolve disagreements with two additional annotators, ultimately establishing the final labeled data as ground truth.
Appendix C Different LLMr
Model Name Quantization Knowledge Cut-off Date
Llama 3.2 3B Instruct Turbo FP8 2023/12
Llama 3.1 8B Instruct Turbo FP8 2023/12
Llama 3.1 70B Instruct Turbo FP8 2023/12
Llama 3.3 70B Instruct Turbo FP8 2023/12
Mistral (7B) Instruct v0.3 FP16 2024/5
gpt-3.5-turbo   — 2021/9
Table 5: Different LLMr to be evaluated. In ScopeQA-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cut-off date 5/22/2024.
Appendix D Out-of-scope Example
Document:
Justin Jefferson, CeeDee Lamb, NFL Injury Statuses and Fantasy Impact for Week 3 Stephen Maturen/Getty Images
Two of the best wide receivers in the NFL are expected to be on the field for their respective Week 3 games. Justin Jefferson and CeeDee Lamb each received positive prognosis about their injury issues, which is something that can’t be said for the rest of the stars across the NFL. Christian McCaffrey is already on injured reserve, Deebo Samuel is out for a weeks and a slew of other running backs and wide receivers are dealing with ailments that could keep them out of Week 3. Below is a look at all of the significant injuries that could affect fantasy football matchups across Week 3. Justin Jefferson Off Injury Report
Justin Jefferson was taken off the Minnesota Vikings injury report on Friday. Jefferson’s status was up in the air because of a quad injury, but he practiced well enough this week that the injury is not a concern. The superstar wide out will be needed for Minnesota’s home clash with the Houston Texans, which has the potential to be a high-scoring affair. Jefferson is always the primary target in Minnesota when healthy, but he should have more targets in Week 3 because Jordan Addison and T.J. Hockenson are still out. Jefferson earned seven targets from Sam Darnold in Week 2. Only running back Aaron Jones had more than four targets against the New York Giants. Houston’s defense may give Jefferson some fits, led by cornerback Derek Stingley Jr. but it has allowed 330 receiving yards to opposing wide outs on just 20 catches through two weeks. Jefferson is an automatic start whenever he’s healthy, and his star power may be needed more on certain fantasy rosters in Week 3 depending on how many injuries affect a single roster.
Question: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
Explanation: This question is out-of-scope as it asks about "Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft" which was never discussed in the document.
Appendix E Statistics of ScopeQA
Attr Gold Silver
# of documents 45 1955
# of in-scope questions 103 8,949
# of out-of-scope questions 113 10,105
average words per document 280 276
average words per in-scope question 16 16
average words per out-of-scope question 16 16
Table 6: Statistics of ScopeQA. Gold is sampled from crawled 2000 news and annotated by annotators (Appendix B).
Appendix F Prompt
F.1 Extract Claims Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and list {num_fact} most important facts it contains. Each fact should be stated in a clear, standalone sentence with sufficient context to be understood independently, avoiding undefined pronouns. Ensure that each fact is directly derived from the document and does not include any information not mentioned within it.
Document:
"""{document}"""
{num_fact} most important facts:
F.2 Recover Missing Claims Prompt
Read the document below with a list of {num_fact} facts it contains. Note that some of the facts are missing represented by (missing). Your task is to guess the missing facts could have said and complete the missing facts. Each fact should be stated in a clear, standalone sentence with sufficient context to be understood independently, avoiding undefined pronouns. Please complete the missing facts and return all the {num_fact} facts in original order. You must return all the {num_fact} facts.
Document:
"""document"""
Completed list of facts:
F.3 Out-of-scope Judgement Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and the question. Check if the question contains any confusing parts: mentions of people, locations, organizations, or other entities that are not present in the document
Document:
"""{document}"""
Question:
{question}
Does the question contain any confusing parts: mentions of people, locations, organizations, or other entities that are not present in the document? Please start by explaining your reasoning. End your response with ’The answer is: Yes or No.’:
F.4 Response Defusion Judgement Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions. Read the document and answer the question based on the document.
Document:
"""{example_document}"""
Question:
{example_questions}
Answer:
{example_LLM_responses}
{Defusion Instruction (Appendix  F.5)}
{example_defuse_reasoning}
Read the document and answer the question based on the document.
Document:
"""{document}"""
Question:
{question}
Answer:
{LLM_response}
{Defusion Instruction (Appendix  F.5)}
F.5 Defusion Instruction
The question presented is confusing and unanswerable based on the document provided. Read the response carefully. If the response attempts to answer the question, like answering ’Yes’ to agree with the question or ’No’ to disagree with the question, you should immediately conclude with ‘The answer is: No.’ even if it later acknowledges the question’s confusion. Otherwise, determine if the response clarifies that the question cannot be answered or if the response indicates that the information needed to answer the question is not available or not mentioned in the document. Begin with your reasoning and finally conclude with ‘The answer is: Yes.’ or ‘The answer is: No.’. Note that if the response attempts to answer the question at any point, like answering ’Yes’ to agree with the question or ’No’ to disagree with the question, you should immediately conclude with ‘The answer is: No.’ even if it later acknowledges the question’s confusion:
F.6 RAG Basic
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
Document:
"""{document}"""
Question:
{question}
Answer:
F.7 RAG Two-shot
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
Document:
"""{example_document}"""
Question:
{two_example_questions}
Answer:
{two_example_responses}
Read the document and answer the question based on the document.
Document:
"""{document}"""
Question:
{question}
Answer:
F.8 RAG Zero-shot-CoT
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and reason step by step to answer the question based on the document. If the question cannot be answered using the document, state explicitly that the question cannot be answered.
Document:
"""{document}"""
Question:
{question}
Answer:
F.9 In-scope Question Generation Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions. Read the document attentively and compile a numbered list of the top {num_q} questions that the document directly answers. Ensure each question is clear, accurate, and devoid of confusion, false assumptions, undefined pronouns, or misinformation. Avoid referencing people, locations, organizations, or other entities not explicitly mentioned in the document. Construct each question to be thought-provoking, containing between 13 to 18 words, and sufficiently detailed to avoid being overly straightforward.
Document:
"""{document}"""
Questions:
F.10 Out-of-scope Question Generation Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and review the list of hallucinated facts. For each hallucinated fact, craft a single, specific and concise question containing 13 to 18 words that incorporate the key element of the fact, ensuring the question is intentionally confusing. The question should not be answerable using any information present in the document. The question should not combine multiple queries and each question should address only one specific aspect. If a question cannot be formulated for a particular hallucinated fact, you may omit it.
Document:
"""{document}"""
hallucinated facts:
{hallucinated_facts}
Questions:
F.11 Remove Claims Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document below with a list of {num_true_fact} ground-truth facts it contains and a list of {num_false_fact} hallucinated facts that are not supported by the document. Your task is to remove any hallucinated facts that can be supported by either the document or the {num_true_fact} ground-truth facts. Please only return the remaining hallucinated facts, along with their original order numbers.
Document:
"""{example_document}"""
{num_true_fact} ground-truth facts:
{true_facts}
{num_false_fact} hallucinated facts:
{false_facts}
Remaining hallucinated facts:
{remained_facts}
Read the document below with a list of {num_true_fact} ground-truth facts it contains and a list of {num_false_fact} hallucinated facts that are not supported by the document. Your task is to remove any hallucinated facts that can be supported by either the document or the {num_true_fact} ground-truth facts. Please only return the remaining hallucinated facts, along with their original order numbers.
Document:
"""{document}"""
{num_true_fact} ground-truth facts
{true_facts}
{num_false_fact} hallucinated facts:
{hallucinated_facts}
Remaining hallucinated facts:
Appendix G Expenses
Generating the ScopeQA dataset using GPT-4o-mini costs approximately \$80. For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/. Excluding the dataset generation, all other experiments cost around \$100.
Appendix H Embedding Classifier
H.1 Implementation Details
We train the classifier on ScopeQA-Silver, which is split as: 80% for training, 10% for evaluation, and 10% for testing. After training, we test the classifier on ScopeQA-Gold data.
We append the unused-token ID at the end of the tokenized input sequence. After the LLM processes the input, we extract the unused-token’s hidden state from the model’s last hidden layer output. Since this token is not part of the LLM’s training and does not carry predefined meanings, it serves as a placeholder to aggregate information from the model’s attention over all preceding tokens in the sequence. As the final token in the input, its embedding encapsulates the model’s representation of the entire input sequence, reflecting the information inferred from the input sequence.
The unused-token’s hidden state is used as input features for a 2-layer MLP, which is optimized with a BCE loss. A dropout of 0.1 is applied after the first layer. We use Adam optimizer with 1e-4 learning rate. We train each classifier for 10 epochs with batch size 8. We save the best validation checkpoint and report its performance on ScopeQA-Gold.
H.2 Unused-Token
For the Llama family, we use the <|reserved_special_token_0|> as the unused-token. This token is not trained and is reserved for future use cases such that the users do not need to resize the vocabulary. For Mistral-7B-v0.3, we use [control_555] as the unused-token because Mistral’s official documentation indicated that they reserve 768 control tokens for future use, and none of them are used during training. 555 is a random choice.
H.3 Sentence BERT NLI baseline
To create a baseline classification performance, we adapt NLI style training using an off-the-shelf sentence-transformers/all-mpnet-base-v2 model from HuggingFace. Specifically, we follow Reimers and Gurevych (2019) to extract document and question [CLS] embeddings $u$ and $v$ to use $[u;v;|u-v|]$ as the input feature for the 2-layer MLP to classify.