### v3

# ELOQ: Resources for Enhancing LLM Detection of Out-of-Scope Questions

Abstract.
Large Language Models (LLMs) are widely used in Conversational AI systems to generate responses to user inquiries. However, many natural questions lack well-defined answers. While existing studies primarily focus on question types such as false premises, they often overlook out-of-scope questions, where the provided document is semantically highly similar to the query but does not contain the required answer. In this paper, we propose a guided hallucination-based method to efficiently generate a diverse set of out-of-scope questions from a given document corpus. We then evaluate multiple LLMs based on their effectiveness in confusion detection and appropriate response generation. Furthermore, we introduce an improved method for detecting such out-of-scope questions, enhancing the reliability of LLM-based question-answering systems.
Large Language Models, Question Answering, Out-of-Scope Question
*Both authors contributed equally to this work.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processing
1. Introduction
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents (Lewis et al., 2020; Guu et al., 2020; Gao et al., 2024). Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and provides it to a large language model (LLM) that generates a response. RAG performance is evaluated on dimensions such as response accuracy, faithfulness, completeness, answer and context relevance, and style (Es et al., 2024; Saad-Falcon et al., 2024; Liu et al., 2024; Zheng et al., 2023). Errors may stem from a document-to-question mismatch, hallucinated parts of the response, or misunderstanding of the document or the question.
Many user questions have no direct answer: some rely on a false premise, others are ambiguous in their intent, and some are out-of-scope for the knowledge base. Prior studies indicate that about 25% of natural questions contain false assumptions (Yu et al., 2023) and over 50% are ambiguous (Min et al., 2020a). Additionally, out-of-scope errors have been identified as the leading cause of enterprise AI assistants generating responses that appear convincing but are, in fact, incorrect (Maharaj et al., 2024). A well-designed RAG system should implement an exception-handling mechanism that detect confusing questions and responds by clarifying the confusion, retrieving additional documents, or seeking assistance.
Dataset Source Category Cutoff knowledge Annotation Method
AMBIGQA (Min et al., 2020a) Wiki Multi-answers Human
CREPE (Yu et al., 2023) Reddit False Premise Human
FalseQA (Hu et al., 2023)  False Premise Human
KG-FPQ (Zhu et al., 2024) Wiki False Premise LLM
FAITH (Yuan et al., 2024) Wiki False Premise LLM
FLUB (Li et al., 2024a) Baidu Tieba Cunning Texts Human
ELOQ News Out-of-scope Human+LLM
Table 1. ELOQ vs Existing Datasets
Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type. As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge, but often neglects out-of-scope queries for which no answers can be found in retrieved documents. Out-of-scope questions can mislead LLMs into generating hallucinated answers that appear correct, but are incorrect, which represents the most dangerous type of error (Maharaj et al., 2024). Our focus is to mitigate this type of confusion.
This paper proposes a framework for automatically generating out-of-scope questions and evaluates its effectiveness in training classifiers to detect out-of-scope situations. While queries in MS MARCO (Nguyen et al., 2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cutoff. ELOQ fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to reduce human annotation effort. Our contributions are as follows:
  • We propose and implement an efficient method for generating synthetic out-of-scope questions from a given set of documents.
  • We demonstrate the utility of our synthetic data in training out-of-scope exception detectors.
  • We conduct a comparative evaluation of multiple LLMs as RAG agents to measure their confusion detection accuracy.
  • Our code  [^1] https://github.com/zhiyuanpeng/ELOQ.git and dataset  [^2] https://huggingface.co/datasets/zhiyuanpeng/ELOQ are made publicly available.
2. Related Work
2.1. Benchmark Datasets
Recent work on LLM responses to confusing questions primarily focuses on stand-alone questions answered using the LLM’s pretraining knowledge, without a supporting context document. These questions may be naturally collected from online sources (Li et al., 2024b; Yu et al., 2023), written by human annotators (Hu et al., 2023), or synthetically generated (Zhu et al., 2024; Yuan et al., 2024). Released benchmark datasets with confusing questions include KG-FPQ (Zhu et al., 2024), FalseQA (Hu et al., 2023), FLUB (Li et al., 2024b), and CREPE (Yu et al., 2023). Earlier research on ambiguous open-domain questions (Min et al., 2020b) introduced AmbigNQ, a dataset of naturally ambiguous questions. Table 1 presents a comparison of these benchmark datasets.
2.2. Question Generation
(Zhu et al., 2024; Yuan et al., 2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of (subject, relation, object) using Wikipedia data, replacing the object with a similar but incorrect one, and then filling in a template to generate a false premise question. Content-grounded (non-confusing) synthetic data generation has been studied in (Yehudai et al., 2024; Wang et al., 2023a; Lee et al., 2023). Given the grounding passages, these methods generate data entries using few-shot prompting with task-specific examples, followed by filtering to ensure data quality, faithfulness, and diversity.
2.3. Mitigation
Mitigation is commonly achieved through in-context learning  (Zhu et al., 2024; Li et al., 2024b), chain-of-thought (CoT)  (Wei et al., 2022) and fine-tuning. Experiments show that fine-tuning outperforms both in-context learning and CoT reasoning (Hu et al., 2023). CoT reasoning leads to inconsistent improvements, whereas in-context learning improves performance as the number of examples increase (Li et al., 2024b). Yuan et al. (2024) identifies approximately 1% of attention heads as responsible for confusion caused by false-premise questions and demonstrates a 20% performance gain by constraining these heads.  Kim et al. (2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.
Algorithm 1 The guided hallucination method
1:Call LLMq to extract from $d$ a list of $n$ claims $c_{1},\ldots,c_{n}$; batching claims reduces LLMq-calls (Appendix B.1)
2:Partition $\{c_{1},\ldots,c_{n}\}$ into disjoint subsets $S_{1},\ldots,S_{k}$, e.g. $S_{j}=\{c_{i}:i\mathop{\textrm{mod}}3=j-1\}$ for $j=1,2,3$
3:for several rounds (e.g. 3), and in each round, for all $S_{j}$ in partition $(S_{1},\ldots,S_{k})$ do
4:    In the claims list $c_{1},\ldots,c_{n}$ replace all claims $c_{i}$ where $i\in S_{j}$ by text “(missing)
5:    Provide LLMq with this modified list (but not $d$) and prompt it to recover the missing claims (Appendix B.2)
6:    In the claims list $c_{1},\ldots,c_{n}$, replace the missing claims with the claims recovered in the above step.
7:end for
8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only out-of-scope claims (Appendix B.11)
9:Call LLMq to generate one short question per each novel claim, focusing on one key element of it (Appendix B.10)
10:Call LLMq to filter out any questions answerable in the context of $d$ (Appendix B.3)
3. Methods
We developed the ELOQ data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions for a given document. We assume that the user has limited knowledge of the document, resulting in asking questions that appear relevant but are actually out-of-scope and cannot be answered based on the document’s content. In this step, given a standard prompt $p$, an out-of-scope question $q$, and a relevant document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question – a process we refer to as defusion (or de-confusion) – rather than attempting to answer the question directly and risking hallucination.
Algorithm 2 Evaluation steps
1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix B.6,  B.7,  B.8)
2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix B.3)
3:Aggregate $m$ predictions by majority vote
4:if $q$ is generated as out-of-scope then
5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix B.4)
6:     Aggregate $m$ predictions by majority vote
7:end if
Attr Gold Silver
# of documents 45 1955
# of in-scope questions 103 8,949
# of out-of-scope questions 113 10,105
average words per document 280 276
average words per in-scope question 16 16
average words per out-of-scope question 16 16
Table 2. Statistics of ELOQ. Gold is sampled from crawled 2000 news and annotated by annotators (Section 3.4).
3.1. Data Collection
We focus on the scenario where the domain knowledge is maintained separately from the LLM in a document database. Hence, we prefer documents that are novel to the evaluated LLMs, such as news articles published after all pretraining cutoff dates, ensuring that LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Instead, they must generate responses grounded in their understanding of the news content. Therefore, using the Newscatcher (Bugara et al., 2020), we collected 200 news articles published between Jan 1, 2024, and September 30, 2024, for each topic333sport, business, science, food, politics, travel, entertainment, music, news, tech. We require each document to be short enough to fit in the LLM prompt along with instructions and examples but long enough to make at least 4 to 10 separate claims (i.e. a few paragraphs long). Thus, we select news articles with more than 150 words and, for each, sequentially extract sentences from the beginning until the word count exceeds 300. Finally, we collected 2,000 news and sampled a subset for human annotation. We represent the human-annotated data as Gold and the remaining data as Silver. Table 2 presents the dataset statistics.
3.2. Question Generation
We introduce our guided hallucination method, designed to generate out-of-scope questions from a given document. The process consists of three main steps: claim extraction, hallucination injection, and question generation.
3.2.1. Claim Extraction
The first step involves extracting claims from the document $d$ to serve as the basis for generating out-of-scope questions. Since our goal is to generate questions that appear related to $d$ but are actually unanswerable using its content, we start by obtaining a structured set of claims from the document. To achieve this, we use a LLM denoted as LLMq, to extract a set of $n$ claims $c_{1},c_{2},…,c_{n}$ from $d$. These claims represent key factual statements or assertions made in the document. We extract the claims in batches to minimize the number of LLM calls and improve computational efficiency. Once the claims are extracted, we partition them into disjoint subsets $S_{1},S_{2},…,S_{k}$. A simple partitioning scheme is used, such as:
$S_{j}=\{c_{i}\mid i\mod 3=j-1\},\quad\text{for }j=1,2,3.$
This ensures that claims are distributed evenly among the subsets, facilitating an efficient iterative process in the next step. Partitioning the claims allows us to selectively manipulate certain claims while preserving the overall document structure.
3.2.2. Hallucination Injection
The second step is to generate hallucinated claims that are similar to the factual claims generated in Section 3.2.1 but contain altered or missing information. The process follows an iterative masking and recovery approach over multiple rounds (e.g. 3). For each round:
  • For all claims in one subset $S_{j}$, we replace them with placeholders “(missing)”.
  • The modified claims list (with missing claims) is then provided to LLMq without access to $d$. The model is prompted to recover the missing claims based on the remaining claims in the list.
  • The recovered claims are inserted back into the claims list, replacing the placeholders.
  • This process is repeated for all subsets $S_{1},…,S_{k}$ across several rounds.
Finally, we discard claims that are explicitly supported by $d$. This ensures that only truly out-of-scope claims remain in our dataset.
3.2.3. Question Generation
The final step converts the hallucinated claims into out-of-scope questions while ensuring they cannot be answered using the document. Specifically, for each hallucinated claim, LLMq generates a concise and precise question focusing on a key aspect of the claim. The question is designed to appear related to the document while being unanswerable based on its content. Once the questions are generated, we further filter them by ensuring they are truly unanswerable. This is achieved by prompting LLMq to verify whether the generated questions can be answered using $d$. Any question that remains answerable in the context of $d$ is discarded.
In addition to generating out-of-scope questions, we also prompt LLMq to generate in-scope questions for comparison. The complete question generation process, along with its prompts, is outlined in Algorithm 1.
Group Name Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
Out-of-Scope defusion Out-of-Scope Defusion Out-of-Scope Defusion Out-of-Scope Defusion
1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
  — 87.50 94.59
2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
  — 91.67 90.00
3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
  — 98.61 100.00
Table 3. Agreement scores and the performance of our method on the ELOQ-Gold.
3.3. Evaluation
We first assessed the quality of ELOQ by sampling a subset of questions (ELOQ-Gold) and manually verifying whether they were truly out-of-scope (Section 3.4). Next, we evaluated how well various LLMr models could defuse the out-of-scope questions in ELOQ. Given the large number of out-of-scope questions, manual verification of LLMr responses is not feasible. Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question. To validate AutoDefuseEval, we compared its performance with human annotations (ELOQ-Gold) on a sampled dataset (Section 3.4). Formally, we define our tasks as follows:
Question generation::
What is the quality of out-of-scope and in-scope questions generated by Algorithm 1? (Section 3.4)
Defusion detection::
How accurately can AutoDefuseEval detect whether an LLMr response to a out-of-scope question defuses the confusion? (Section 3.4)
Out-of-scope response::
How often does an LLMr successfully defuse an out-of-scope question using different prompting methods? (Section 4.1)
Out-of-scope detection::
How accurately can LLMs identify which context-grounded questions are out-of-scope and require special handling? (Section 4.2)
Algorithm 2 has the steps we run for evaluating the above tasks. We run steps 1, 2, and 3 for “out-of-scope detection” and 1, 4, 5, 6, and 7 for “out-of-scope response”. Following the self-consistency method (Wang et al., 2023b), we perform multiple LLM calls in steps  2 and  5 of Algorithm 2, taking the majority vote to determine the final label.
3.4. Human Annotation
Figure 1. Confusion matrix of out-of-scope and defusion on ELOQ-Gold.
Following the methodology from (Cochran, 1977), we sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators. As shown in Table 3, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is out-of-scope and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency. Two additional annotators manually resolved all disagreements to establish ground truth labels. All annotators were computer science graduate students recruited via email and incentivized with complimentary meals. They were trained by first reviewing the annotation guidelines and then labeling a small dataset previously annotated by the authors. The annotation guidelines are released along with the data and code.
We used two metrics to assess our method: “Annotator Acc” and “Group’s Agree Acc.” “Annotator Acc” treats each annotator’s labels as ground truth, computing the accuracy of the generated labels. “Group’s Agree Acc” uses the instances where both annotators within the same group agreed, treating this consensus as the ground truth. This method yielded higher accuracy than “Annotator Acc,” leading us to resolve disagreements with two additional annotators, ultimately establishing the final labeled data as ground truth.
As reflected in “Ground Truth Acc,” in Table  3, our method achieved 94.91% accuracy in generating out-of-scope and in-scope questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion. As shown in Figure 1, the confusion matrix on the left aligns closely with the ground truth, with minimal errors (5 false negatives and 6 false positives). The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
Model Name Quantization Knowledge Cutoff Date
Llama 3.2 3B Instruct Turbo FP8 2023/12
Llama 3.1 8B Instruct Turbo FP8 2023/12
Llama 3.1 70B Instruct Turbo FP8 2023/12
Llama 3.3 70B Instruct Turbo FP8 2023/12
Mistral (7B) Instruct v0.3 FP16 2024/5
gpt-3.5-turbo   — 2021/9
Table 4. Different LLMr being evaluated. In ELOQ-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cutoff date of 5/22/2024.
4. Experimental Results
Prompt Model business entm food music news politics science sport tech travel Avg Std Dev
Basic GPT-3.5 18.06 10.96 12.32 14.27 13.59 15.95 10.15 $18.37^{\star}$ 15.59 9.63 13.89 2.97
Llama 3.2 3B 63.59 61.55 61.09 62.07 63.49 64.01 53.35 $71.71^{\star}$ 63.69 56.51 62.11 4.60
Mistral 7B v0.3 68.40 67.63 60.88 67.74 67.96 66.86 56.07 $76.23^{\star}$ 66.87 61.06 65.97 5.20
Llama 3.1 8B 71.25 69.42 67.56 72.04 67.16 71.89 59.73 $77.90^{\star}$ 70.15 63.49 69.06 4.74
Llama 3.1 70B 71.54 70.22 67.56 71.07 67.66 70.75 61.51 $79.76^{\star}$ 70.67 68.04 69.88 4.33
Llama 3.3 70B 71.25 72.71 68.17 73.51 69.84 69.90 60.04 $\textbf{82.42}^{\star}$ 73.23 68.78 70.98 5.29
Two-shot GPT-3.5 65.65 65.24 57.60 65.88 62.20 66.48 52.41 $71.71^{\star}$ 59.90 56.83 62.39 5.43
Llama 3.2 3B 80.77 79.28 76.08 79.47 76.79 78.25 72.28 $84.09^{\star}$ 79.90 71.64 77.85 3.61
Mistral 7B v0.3 80.86 77.69 74.54 78.79 79.17 79.11 67.47 $86.84^{\star}$ 78.15 71.43 77.40 5.02
Llama 3.1 8B 83.32 80.98 80.18 84.16 78.77 82.53 72.70 $87.92^{\star}$ 82.77 78.41 81.17 3.87
Llama 3.1 70B 85.18 83.37 80.60 85.83 79.76 82.24 75.84 $\textbf{90.57}^{\star}$ 83.08 79.79 82.62 3.84
Llama 3.3 70B 85.48 81.27 80.39 84.26 79.27 82.34 75.84 $88.90^{\star}$ 83.49 78.62 81.99 3.57
Zero-shot-CoT GPT-3.5 63.00 61.95 55.95 64.03 62.40 62.58 53.03 $71.71^{\star}$ 59.69 56.30 61.07 4.96
Llama 3.2 3B 83.91 78.09 77.21 79.96 78.77 81.77 73.85 $85.46^{\star}$ 80.41 75.77 79.52 3.39
Mistral 7B v0.3 77.72 78.39 70.94 78.59 79.07 77.02 67.26 $84.77^{\star}$ 76.41 72.06 76.22 4.69
Llama 3.1 8B 79.39 78.09 74.85 78.40 78.67 79.20 70.29 $83.89^{\star}$ 77.64 74.50 77.49 3.43
Llama 3.1 70B 90.28 87.85 87.27 88.56 86.51 89.08 83.68 $91.55^{\star}$ 86.97 85.19 87.69 2.21
Llama 3.3 70B 93.52 91.24 90.55 92.57 90.77 92.12 89.12 $\textbf{95.09}^{\star}$ 93.85 90.90 91.97 1.71
Table 5. Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ELOQ-Silver dataset across diverse news topics. For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics within the same prompting method. Bold values are the maximum values for each topic across all the LLMs within the same prompting method. “entm” is the abbreviation of “entertainment”.
4.1. Out-of-Scope Response
In this section, we examine LLMs’ ability to defuse out-of-scope questions. The selected LLMs are listed in Table 4. Most of these models have a knowledge cutoff date of December 2023, which predates the publication of the news articles in ELOQ. As a result, they must generate responses based solely on their understanding of the provided news content. As shown in Table 5, we evaluated LLMs on their ability to defuse out-of-scope questions with three prompting methods (Appendix B.6,  B.7,  B.8). As shown in Table 5, a) Two-shot prompt boosts all evaluated LLMs’ accuracy substantially, especially GPT-3.5, likely due to the examples clarifying how to respond to out-of-scope questions. In contrast, Zero-shot-CoT (Kojima et al., 2022) merely invites the LLM to “reason step by step” beats Two-shot on three out of six LLMs. b) Larger models (70B) benefit more from Zero-shot-CoT than from Two-shot prompting, achieving accuracy gains of 5.07% to 9.98%, whereas smaller models (3B to 8B) show only marginal changes, ranging from -3.68% to 1.67%. This difference likely stems from larger models’ ability to better utilize their reasoning capabilities and extensive knowledge; c) Most LLMs perform worst on the “science” topic and best on “sport”. This is expected, as scientific content tends to contain more implicit premises assumed by the author rather than explicitly stated facts, thus requiring more domain knowledge. In contrast, sports content is simpler and unambiguous, with facts often clearly stated in the document.
4.2. Out-of-Scope Detection
The main challenge in preventing an LLM from hallucinating or generating incorrect answers to out-of-scope questions is detecting such scenarios. Once we can accurately identify these cases, we can handle them easily by using a prompt and in-context examples specifically designed for out-of-scope questions. We use two methods for out-of-scope detection. First is a prompt-based method where we simply prompt the LLM to do such detection (Appendix B.3), denoted as “Direct Generate” in Figure 2. Second is a classification method where we train a binary classifier on ELOQ-Silver using the embeddings of an appended unused-token as features, denoted as “Unused-token Classification” in Figure 2.
Figure 2. Evaluation of out-of-scope detection on ELOQ-Gold.
4.2.1. Unused-token Classification
The out-of-scope detection problem can be framed simply as a binary classification task, where we determine if a given query $q$ is out-of-scope or not for a given document $d$. The binary label is $y\in\{0,1\}$ where $y=1$ indicates the query is relevant to the document and $y=0$ indicates the query is out-of-scope. The objective is to learn a classification function: $f:(q,d)\mapsto p(y=1\mid q,d)$, which estimates the likelihood of the query-document pair to be in-scope. We would like to test whether the LLM’s internal features contain enough information to classify a query-document pair effectively to simulate the effect that the LLM “realizes” the out-of-scope scenario before generating a response. To achieve this, we use the sequence representation of the entire input – denoted as $x=\text{concat}(I,q,d,)$ where $I$ is the instruction prompt provided to the LLM (e.g. “Can this document answer the question?”). We extract the sequence representation by appending a reserved unused-token to the end of the input sequence and extracting its representation from the final hidden state of the last transformer layer. The rationale behind using unused-token is that, since LLM developers reserve these tokens for potential future use in specialized training tasks, they are untrained and uninfluenced by any prior data. Thus, the unused-token does not have any predefined meanings and can be viewed as a blank buffer, aggregating interactions among all the input tokens and capturing an uninfluenced representation of the sequence. We also tried using the <eos> token, which yielded similar results.
The classifier is a two-layer MLP that takes the unused-token representation as input and is optimized against the binary label $y$. This approach is commonly known as “probing” (Ye et al., 2024), where the LLM weights are frozen, and the classifier can be viewed as an external probe to assess the information contained in the model’s internal representations. The performance reported in Figure 2 was tested on ELOQ-Gold, the human-labeled data. For more details about the training, please refer to Appendix D. We also provide a “Sentence BERT NLI” baseline where we adopt NLI style training using the off-the-shelf  [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 model from HuggingFace. Specifically, we follow (Reimers and Gurevych, 2019) to extract document and question [CLS] embeddings $e_{d}$ and $e_{q}$ to use $[e_{q};e_{q};|e_{d}-e_{q}|]$ as the input features.
4.2.2. Results Discussion
For the “Direct Generation” approach, we can observe a clear positive correlation between model size and accuracy, aligning with the general trend of increased capability with larger models. However, due to the proprietary nature of OpenAI’s models and computational constraints, we are only able to obtain hidden states to train the binary classifier for three LLMs: Llama 3.2 3B, Mistral 7B, and Llama 3.1 8B. Notably, the classifiers consistently outperform the “Direct Generation” method, indicating that the LLM’s internal representations encode the necessary information for out-of-scope detection, yet the model may fail when directly prompted. Our dataset, ELOQ, demonstrates its usefulness by enabling even smaller models – such as Llama 3.2 3B-based classifier – to achieve accuracy comparable to significantly larger models that rely on direct generation, including Llama 3.1 70B and GPT-4o-mini. Moreover, as shown in Figure 2, all models and methods outperform the “SBERT NLI” baseline, suggesting that out-of-scope detection is non-trivial and requires a deeper and more complex understanding of linguistic relationships that larger models capture during pretraining and fine-tuning. Lastly, in our exploratory experiments, we trained a classifier on GPT-2, but its accuracy is no better than random guessing, further highlighting the importance of model capacity in leveraging internal representations for classification.
Model Recall@1 Recall@5 Recall@10 MRR
BM25 (Pérez-Iglesias et al., 2009) 0.4637 0.6880 0.7487 0.5653
SBERT (Reimers and Gurevych, 2019) 0.4685 0.7058 0.7858 0.5773
BGE (Xiao et al., 2023) 0.5326 0.7604 0.8267 0.6352
Stella (Zhang and FulongWang, 2024) 0.4875 0.7376 0.8128 0.5994
Linq (Kim et al., 2024) 0.5603 0.7890 0.8492 0.6618
Table 6. Retrieval performance of out-of-scope questions on ELOQ
4.3. Semantic Relevance
To demonstrate that the generated out-of-scope questions are indeed semantically similar to the document they are generated from, we conduct retrieval experiments using all out-of-scope questions against our document corpus (2000 document). We employ the following retrieval models: BM25  [^5] https://github.com/castorini/pyserini, SBERT  [^6] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, bge-large-en-v1.5  [^7] https://huggingface.co/BAAI/bge-large-en-v1.5, stella_en_1.5B_v5888https://huggingface.co/NovaSearch/stella_en_1.5B_v5, and Linq-Embed-Mistral [^9] https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral. Among these models, BM25 and SBERT are well-known retrievers that serve as established baselines, bge-large-en-v1.5 was accepted in SIGIR 2024 resource track as a high-performance embedding model, stella_en_1.5B_v5 ranks #1 on MTEB’s English v1 dataset (Muennighoff et al., 2022), and Linq-Embed-Mistral leads multiple benchmarks, including BEIR (Kamalloo et al., 2023) and MTEB’s English v2 dataset at the time of writing. As shown in Table 6, the Recall@1 values across all models suggest that for approximately half of the out-of-scope questions, the top-retrieved document is the same one from which the question was generated. When we relax this condition to Recall@10, around 80% of out-of-scope questions retrieve their associated document within the top 10 results. This demonstrates the strong semantic connection between out-of-scope questions and their source documents despite the fact that the document lacks the information needed to answer the question.
Module Accuracy
Base (gpt-4o-mini) 87.61
+ examples 97.35
+ self-consistency (m=3) 98.23
+ self-consistency (m=9) 98.23
Table 7. Evaluation of different prompts for “defusion detection” on ELOQ-Gold.
4.4. Ablation Study
We examined the impact of various prompts on “defusion detection” task. As seen in Table 7, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $m=3$ further boosted performance to 98.23%. However, further increasing $m$ did not lead to additional improvements. Since ELOQ-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ELOQ-Silver dataset. Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix C).
5. Conclusion and Future Works
We introduced ELOQ, a benchmark for enhancing LLM detection of out-of-scope questions for a given document. Human annotation validates that ELOQ achieves 94.91% accuracy in generating out-of-scope and in-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses. We further demonstrated ELOQ’s utility by training out-of-scope detectors. For future work, we plan to apply for OpenAI funding to utilize more powerful models, such as o1, to improve both data quality and scale. Additionally, we aim to explore post-training techniques to fine-tune LLMs on ELOQ, ensuring they align better with human preferences by clarifying the confusing aspects of a question rather than generating hallucinated answers.
Appendix A Out-of-scope Example
Document:
Justin Jefferson, CeeDee Lamb, NFL Injury Statuses and Fantasy Impact for Week 3 Stephen Maturen/Getty Images
Two of the best wide receivers in the NFL are expected to be on the field for their respective Week 3 games. Justin Jefferson and CeeDee Lamb each received positive prognosis about their injury issues, which is something that can’t be said for the rest of the stars across the NFL. Christian McCaffrey is already on injured reserve, Deebo Samuel is out for a weeks and a slew of other running backs and wide receivers are dealing with ailments that could keep them out of Week 3. Below is a look at all of the significant injuries that could affect fantasy football matchups across Week 3. Justin Jefferson Off Injury Report
Justin Jefferson was taken off the Minnesota Vikings injury report on Friday. Jefferson’s status was up in the air because of a quad injury, but he practiced well enough this week that the injury is not a concern. The superstar wide out will be needed for Minnesota’s home clash with the Houston Texans, which has the potential to be a high-scoring affair. Jefferson is always the primary target in Minnesota when healthy, but he should have more targets in Week 3 because Jordan Addison and T.J. Hockenson are still out. Jefferson earned seven targets from Sam Darnold in Week 2. Only running back Aaron Jones had more than four targets against the New York Giants. Houston’s defense may give Jefferson some fits, led by cornerback Derek Stingley Jr. but it has allowed 330 receiving yards to opposing wide outs on just 20 catches through two weeks. Jefferson is an automatic start whenever he’s healthy, and his star power may be needed more on certain fantasy rosters in Week 3 depending on how many injuries affect a single roster.
Question: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
Explanation: This question is out-of-scope as it asks about ”Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft” which was never discussed in the document.
Appendix B Prompt
B.1. Extract Claims Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and list {num_fact} most important facts it contains. Each fact should be stated in a clear, standalone sentence with sufficient context to be understood independently, avoiding undefined pronouns. Ensure that each fact is directly derived from the document and does not include any information not mentioned within it.
Document:
”””{document}”””
{num_fact} most important facts:
B.2. Recover Missing Claims Prompt
Read the document below with a list of {num_fact} facts it contains. Note that some of the facts are missing represented by (missing). Your task is to guess the missing facts could have said and complete the missing facts. Each fact should be stated in a clear, standalone sentence with sufficient context to be understood independently, avoiding undefined pronouns. Please complete the missing facts and return all the {num_fact} facts in original order. You must return all the {num_fact} facts.
Document:
”””document”””
Completed list of facts:
B.3. Out-of-scope Judgement Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and the question. Check if the question contains any confusing parts: mentions of people, locations, organizations, or other entities that are not present in the document
Document:
”””{document}”””
Question:
{question}
Does the question contain any confusing parts: mentions of people, locations, organizations, or other entities that are not present in the document? Please start by explaining your reasoning. End your response with ’The answer is: Yes or No.’:
B.4. Response Defusion Judgement Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions. Read the document and answer the question based on the document.
Document:
”””{example_document}”””
Question:
{example_questions}
Answer:
{example_LLM_responses}
{Defusion Instruction (Appendix  B.5)}
{example_defuse_reasoning}
Read the document and answer the question based on the document.
Document:
”””{document}”””
Question:
{question}
Answer:
{LLM_response}
{Defusion Instruction (Appendix  B.5)}
B.5. Defusion Instruction
The question presented is confusing and unanswerable based on the document provided. Read the response carefully. If the response attempts to answer the question, like answering ’Yes’ to agree with the question or ’No’ to disagree with the question, you should immediately conclude with ‘The answer is: No.’ even if it later acknowledges the question’s confusion. Otherwise, determine if the response clarifies that the question cannot be answered or if the response indicates that the information needed to answer the question is not available or not mentioned in the document. Begin with your reasoning and finally conclude with ‘The answer is: Yes.’ or ‘The answer is: No.’. Note that if the response attempts to answer the question at any point, like answering ’Yes’ to agree with the question or ’No’ to disagree with the question, you should immediately conclude with ‘The answer is: No.’ even if it later acknowledges the question’s confusion:
B.6. RAG Basic
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
Document:
”””{document}”””
Question:
{question}
Answer:
B.7. RAG Two-shot
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
Document:
”””{example_document}”””
Question:
{two_example_questions}
Answer:
{two_example_responses}
Read the document and answer the question based on the document.
Document:
”””{document}”””
Question:
{question}
Answer:
B.8. RAG Zero-shot-CoT
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and reason step by step to answer the question based on the document. If the question cannot be answered using the document, state explicitly that the question cannot be answered.
Document:
”””{document}”””
Question:
{question}
Answer:
B.9. In-scope Question Generation Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions. Read the document attentively and compile a numbered list of the top {num_q} questions that the document directly answers. Ensure each question is clear, accurate, and devoid of confusion, false assumptions, undefined pronouns, or misinformation. Avoid referencing people, locations, organizations, or other entities not explicitly mentioned in the document. Construct each question to be thought-provoking, containing between 13 to 18 words, and sufficiently detailed to avoid being overly straightforward.
Document:
”””{document}”””
Questions:
B.10. Out-of-scope Question Generation Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and review the list of hallucinated facts. For each hallucinated fact, craft a single, specific and concise question containing 13 to 18 words that incorporate the key element of the fact, ensuring the question is intentionally confusing. The question should not be answerable using any information present in the document. The question should not combine multiple queries and each question should address only one specific aspect. If a question cannot be formulated for a particular hallucinated fact, you may omit it.
Document:
”””{document}”””
hallucinated facts:
{hallucinated_facts}
Questions:
B.11. Remove Claims Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document below with a list of {num_true_fact} ground-truth facts it contains and a list of {num_false_fact} hallucinated facts that are not supported by the document. Your task is to remove any hallucinated facts that can be supported by either the document or the {num_true_fact} ground-truth facts. Please only return the remaining hallucinated facts, along with their original order numbers.
Document:
”””{example_document}”””
{num_true_fact} ground-truth facts:
{true_facts}
{num_false_fact} hallucinated facts:
{false_facts}
Remaining hallucinated facts:
{remained_facts}
Read the document below with a list of {num_true_fact} ground-truth facts it contains and a list of {num_false_fact} hallucinated facts that are not supported by the document. Your task is to remove any hallucinated facts that can be supported by either the document or the {num_true_fact} ground-truth facts. Please only return the remaining hallucinated facts, along with their original order numbers.
Document:
”””{document}”””
{num_true_fact} ground-truth facts
{true_facts}
{num_false_fact} hallucinated facts:
{hallucinated_facts}
Remaining hallucinated facts:
Appendix C Expenses
Generating the ELOQ dataset using GPT-4o-mini costs approximately \$80. For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/. Excluding the dataset generation, all other experiments cost around \$100.
Appendix D Embedding Classifier Implementation Details
We train the classifier on ELOQ-Silver, which is split as: 80% for training, 10% for evaluation, and 10% for testing. After training, we test the classifier on ELOQ-Gold data.
For the Llama family, we use <|reserved_special_token_0|> as the unused-token. This token is not trained and is reserved for future use cases such that the users do not need to resize the vocabulary. For Mistral-7B-v0.3, we use [control_555] as the unused-token because Mistral’s official documentation indicated that they reserve 768 control tokens for future use, and none of them are used during training. 555 is a random choice.
The unused-token’s hidden state is used as input features for a 2-layer MLP, which is optimized with a BCE loss. A dropout of 0.1 is applied after the first layer. We use Adam optimizer with 1e-4 learning rate. We train each classifier for 10 epochs with batch size 8. The best validation checkpoint is saved and its performance is reported on ELOQ-Gold.
Acknowledgements.
We would like to express our sincere gratitude to Xiaoxiao Shang, Ethan Lin, Wei Mo, Zhan Shi, Xinze Ye, and Haisong Wang for their invaluable contributions to data labeling.